Suppose you have a hot dog stand, and when you charge $1 per hot dog you sell 75 hot dogs. But when you raise your price to $2 you only sell 50 dogs. Write an equation for the number of hot dogs you sell as a function of the price you charge. Use "H" for number of Hot dogs, and "P" for the price you charge for your hot dogs. You can assume the function is linear.
Answer by stanbon(48568)

Suppose you have a hot dog stand, and when you charge $1 per hot dog you sell 75 hot dogs. But when you raise your price to $2 you only sell 50 dogs. Write an equation for the number of hot dogs you sell as a function of the price you charge. Use "H" for number of Hot dogs, and "P" for the price you charge for your hot dogs. You can assume the function is linear.
---------------------------------------------------
Make a list of dogs sold based on price charged:
H------p
($1----75)
($2----50)
-----------------------
slope = (75-50)/(2-1) = 25
----
intercept = ?
p = mH + b
75 = 25*1 + b
b = 50
----------------
Equation:
p(H) = 25H + 50
============================
Cheers,
Stan H.